# Also at ] The European Space Agency Centre for Earth Observation, Frascati, Italy

### v3

@@ Abstract
We identify a thermodynamic phase transition in chiral active matter. Low-frequency disorder triggers global synchronisation and energy dissipation, while high disorder activates a topological heat pump, generating an inverse energy cascade. This drives the system towards an Onsager dipole, which can be arrested into a metastable vortex glass if dispersion is insufficient to overcome the defect lattice. We propose topological gas dynamics as a universality class governed by the interplay of active disorder and topological sorting, unifying active swarms and classical inviscid fluids.
@@ I Introduction
Active turbulence, the chaotic spatiotemporal dynamics observed in bacterial swarms [1, 2], cytoskeletal extracts [3], and synthetic colloidal flocks [4, 5], presents a challenge to statistical physics. Unlike classical inertial turbulence, which is driven by high Reynolds numbers ($R_{e}$), active turbulence arises in the overdamped limit ($R_{e}\approx 0$), driven by the continuous injection of energy at the microscale [6, 7].
Despite the viscous dominance in such systems, observations frequently reveal large-scale coherent structures reminiscent of inertial fluid flows [8]. This has ignited a debate regarding the universality class of active matter [3]. Current theories, largely built on the hydrodynamics of active nematics, predict steep energy spectra (typically $E(k)\sim k^{-4}$ or $k^{-3}$) dominated by the proliferation of half-integer defects [9, 10]. While such models successfully describe transient pattern formation, they generally fail to predict the robust inverse energy cascades ($k^{-5/3}$) observed in chiral active fluids [11] and 2D quantum turbulence [12].
The discrepancy implies a missing link in the theoretical framework concerning active matter dynamics. While odd viscosity enables unidirectional flow [13, 14] and recent models suggest swarming is entropy-driven [15], the thermodynamic end-state of such systems remains largely unexplored. Specifically, does the breaking of detailed balance (time reversal symmetry breaking [16]) by intrinsic rotation merely sustain a chaotic steady state, or does it drive the system toward a specific macroscopic attractor?
In this letter, we propose that the answer to the conundrum lies in topological gas dynamics. By simulating chiral active matter as an ensemble of $N$ locally phase-coupled oscillators experiencing a Kuramoto-like phase-interaction [17, 18], we demonstrate that chiral active matter can act as a dual-phase system: a microscopic active bath dominated by enstrophy injection ($k^{-3}$), and a macroscopic renormalised fluid element governed by the conservative advection of topological defects. By filtering out the microscopic singularities, we reveal that the simulated system undergoes a “hidden” inverse energy cascade with a Kolmogorov-scaling ($k^{-5/3}$), indistinguishable from Euler (frictionless, inviscid) turbulence. This cascade functions as a thermodynamic pump, driving the system toward an Onsager dipole, a negative-temperature condensate of vorticity [19], thereby unifying the phenomenology of biological swarms with the rigorous statistical mechanics of point vortices.
@@ II Methodology
II.1 The Model
We define the system as a chiral active fluid composed of $N=50,000$ particles, or discrete agents, in a bounded periodic domain $\mathcal{D}\in\mathbb{R}^{2}$. The state of the $i$-th agent is defined by its position $\mathbf{r}_{i}(t)$ and its internal phase $\phi_{i}(t)\in[0,2\pi)$.
The agents do not interact via pairwise collisions (as in granular matter) but through a local mean field. Following Ref. [17], we define the complex order parameter field $Z(\mathbf{r},t)$ as the convolution of the microscopic agent distribution with a finite-range interaction kernel $G$:
@@ II.1 The Model
where $G(|\mathbf{r}-\mathbf{r}^{\prime}|)$ is the Green’s function of the interaction (modeled as a Gaussian with kernel size $\sigma$), $R(\mathbf{r},t)$ is the local coherence (or order parameter, where $R\approx 1$ implies high local synchronisation and $R\approx 0$ implies a phase singularity, or defect core). $\Psi(\mathbf{r},t)$ is the local mean phase. The convolution kernel $G$ is interpreted in the context of chemical gradients or electric fields, which govern the dense active matter interactions (e.g., bacterial swarms or even electron gases). By modifying $G$, we can potentially simulate anisotropic media (like magnetized plasmas) without changing the agent logic.
@@
$\dot{\phi}_{i}=\underbrace{\omega_{i}}_{\text{Driver}}+\underbrace{a_{0}R(\mathbf{r}_{i})\sin(\Psi(\mathbf{r}_{i})-\phi_{i})}_{\text{synchronisation force}}+\underbrace{\sqrt{2D_{\phi}}\eta_{i}(t)}_{\text{Stochastic noise}},$ (2)
where $\omega_{i}$ are natural frequencies drawn from a power-law distribution $P(\omega)\sim\omega^{-\gamma}$, $a_{0}$ is the coupling Strength, and $\eta_{i}$ is a Gaussian white noise term. $R_{i}$, the local order parameter, ensures that synchronisation is triggered by initially random local alignments in phase. $\omega_{i}$ (and the stochastic term $\eta_{i}$) prevent a “ferromagnetic collapse”, in which the system simply synchronises globally ($\phi_{i}=\text{const}$), stopping all motion (or causing a uniform ballistic drift). By forcing the particles to oscillate with an inherent frequency (or attempt to), the $\omega_{i}$ terms act as frustration, constantly injecting vorticity at the micro-scale, which, as we shall demonstrate, feeds an inverse energy cascade.
@@
where $\mathbf{r}_{i}=(x_{i},y_{i})$. The particles are therefore without conventional inertia: $\sum\mathbf{F}\neq m\mathbf{a}$; the system is fully overdamped, in a low Reynolds number limit [22]. In addition, we observe that if the phase field forms a vortex (phase singularity or topological defect), the velocity field must also form a vortex, introducing singularities in the vorticity field $\nabla\times\mathbf{v}$. To define a valid continuum fluid, we introduce the renormalised fluid element (RFE).
A natural definition of the RFE is given by the mean-field kernel size $\sigma$ in our model (Eq. 1),
$\tau=\frac{\sigma}{v_{0}}=\frac{3\text{ grid units}}{0.5\text{ grid units/s}}=6~\text{s}.$
We define the RFE as the center of mass of a particle’s trajectory during $\tau$, effectively course-graining the singularities:
@@
$R_{i}=\text{angle}\left(\frac{1}{\tau}\sum_{t^{\prime}=t-\tau}^{t}\hat{r}_{i}(t^{\prime})\right)\cdot\frac{L}{\pi},$ (4)
with $L$ being the domain half-width, and,
$\hat{r}_{i}(t)=\exp\left(i\cdot\frac{r_{i}(t)}{L}\cdot\pi\right),$ (5)
@@
being the complex phasor locations mapped to a unit circle (necessitated by the periodic boundary conditions). The transformation Eq. (4) constitutes the regularization of the topological defects in $\phi_{i}$, which become physical flow vortices in $\mathbf{r}_{i}$.
@@
The kernel width $\sigma$ defines the correlation length of the local mean field, establishing a natural scale separation. We define the cut-off wavenumber $k_{c}\equiv 2\pi/\sigma$, which, as we shall demonstrate, marks the transition between the microscopic active bath ($k>k_{c}$) and the macroscopic hydrodynamic limit ($k<k_{c}$). The RFE operator effectively acts as a low-pass filter with cut-off $k_{c}$, isolating the collective, fluid modes.
II.2 Theoretical foundation
To bridge the discrete dynamics and the continuum fluid, we define the macroscopic velocity field $\mathbf{u}(\mathbf{r},t)$ as the local average of the particle velocities. Decomposing the particle phase into the local mean phase $\Psi$ and a fluctuation $\delta\phi_{i}$ (that is, $\phi_{i}=\Psi+\delta\phi_{i}$), and assuming symmetric fluctuations (that is, $\langle\sin\delta\phi\rangle\approx 0$), the coarse-grained velocity becomes
$\mathbf{u}=\langle\mathbf{v}_{i}\rangle\approx v_{0}\langle\cos\delta\phi\rangle(\cos\Psi,\sin\Psi),$ (6)
where we identify the local order parameter $R$,
for a peaked (narrow) Gaussian distribution in $\delta\phi$. $R$ measures the degree of of local synchronisation, and so the macroscopic speed is scaled by local synchronisation:
$|\mathbf{u}|=v_{0}R.$ (8)
Thus, while individual particles move at constant speed $v_{0}$, the renormalised fluid element speed is variable and proportional to the local synchronisation. To express $\mathbf{u}$ in terms of the field potential $\Psi$, we identify the flow direction $\mathbf{\hat{n}}=(\cos\Psi,\sin\Psi)$ with the normalized phase gradient:
$\mathbf{u}=v_{0}R\frac{\nabla\Psi}{|\nabla\Psi|}.$ (9)
$\mathbf{u}\approx R\nabla\Psi.$ (10)
This approximation allows us to utilize the machinery of inviscid fluid mechanics in the bulk of the domain, while the factor $R$ naturally regularizes the dynamics at the topological singularities (see Appendix B for additional details).
The continuity equation for $\mathbf{u}$ then reads,
$\frac{\partial\rho}{\partial t}+\nabla\cdot(\rho\mathbf{u})=0.$ (11)
This formulation identifies the system as formally analogous to the shallow water equations [23, 24], where the active density $\rho$ acts as the fluid height. The phase curvature $\nabla^{2}\Psi$ drives the compression of the fluid elements, explicitly linking topological defects to density fluctuations. Consequently, the inviscid Euler behaviour that we observe in the RFE field in the next section represents the low-Mach-number limit of a compressible system, where the rotational, non-compressive wave modes decouple from, and dominate, the irrotational, compressive wave (acoustic) modes at large scales (see Appendix B).
The hydrodynamic closure is completed by the momentum equation. The non-linearity of the phase evolution generates velocity fluctuations, which manifest as an active Reynolds stress $\sigma_{ij}\sim\langle\delta v_{i}\delta v_{j}\rangle$. As we derive in Appendix B, the renormalization of this stress yields the effective inertial advection term $\lambda(R)(\mathbf{u}\cdot\nabla)\mathbf{u}$. Here, the coefficient $\lambda(R)\approx R^{2}$ functions as an emergent inertial mass determined by local synchronisation. This renders the macroscopic momentum transport isomorphic to that of an Euler fluid. Consequently, the motion of topological defects decouples from the overdamped microscopic kinetics, governed instead by the inertial advection of the renormalised superfluid.
@@ II.2 Theoretical foundation
Finally, we identify the topological conservation law that governs the renormalised fluid. As $\Psi(r)$ is the argument of a single-valued complex field $Z$ (Eq. 1), its circulation along any closed loop $\partial\Sigma$ enclosing a defect core is necessarily quantized [25, 26], mirroring the winding number stability in discrete oscillator lattices [27, 17]:
@@
$\oint_{\partial\Sigma}\nabla\Psi\cdot d\mathbf{l}=2\pi m,\quad(m\in\mathbb{Z})$ (13)
implying that $|\nabla\Psi|=m/r$. Substituting this into our macroscopic velocity definition (Eq. 9), and assuming the order parameter $R$ is locally uniform, the fluid velocity induced by a defect scales as:
$|\mathbf{u}|\approx v_{0}R\frac{m}{r}.$ (14)
This confirms that the topological defects of the mean phase act as point vortices with an effective physical circulation $\Gamma\approx 2\pi mv_{0}R$. Consequently, the interaction energy of the ensemble is given by the Kirchhoff-Onsager Hamiltonian [28, 29]:
@@
Crucially, this derivation highlights that the topological charge driving the Hamiltonian dynamics is scaled by the order parameter $R$. If renormalization is suppressed ($R\to 0$), the effective circulation vanishes ($\Gamma_{eff}\to 0$), causing the Hamiltonian interaction to collapse and the system to arrest.
II.3 The Onsager attractor
In a bounded domain $\mathcal{D}$, the phase space volume $\Omega(E)$ (the number of ways to arrange vortices with total energy $E$) eventually decreases as the energy increases. In the low-energy limit, vortices are paired (dipoles), screening each other. In the high-energy limit, like-signed vortices must cluster together to maximize the distance between positive and negative clumps. The result is that $\frac{dS}{dE}<0\implies T<0$ [19]. Frustration $\omega_{i}$ activates a topological heat pump (injecting enstrophy), pumping the system up the energy ladder into the negative-temperature Onsager regime, consistent with entropy-driven ordering observed in active swarms [15].
II.4 Spectral scaling
@@
which sums contributions of all discrete wave vectors $\mathbf{q}=(q_{x},q_{y})$ that fall within the annular shell where the magnitude $|\mathbf{q}|\approx k$, to the real part of the Fourier transform of the advective transport term, where $\hat{\mathbf{v}}_{\mathbf{q}}^{*}$ is the complex conjugate of the velocity field in Fourier space at wave vector $\mathbf{q}$. A negative plateau in $\Pi(k)$ would conclusively demonstrate an inverse cascade.
Figure 1: The emergence of the Onsager dipole in chiral active matter simulation, shown by the simulation’s evolution through seven temporal snapshots, with top row representing the raw particle point-clouds while the bottom row plots topological charge (winding number, Eq. 13) with a colorscale.
Figure 2: Energy characteristics of the simulation run in Figure 1. Panel a) shows the defect (phase singularity) count, with a $t^{-0.75}$ powerlaw scaling [36] with a dashed, green line. Panel b) shows the spectral flux (Eq. 16) for the RFE (blue, left $y$-axis) and raw (red, right $y$-axis) fields, taking the median of the entire simulation. Panel c) shows the energy spectrum $E(k)$ for the RFE (blue) and raw (red) fields; a spectral scaling of $k^{-5/3}$ and $k^{-3}$ is indicated with dotted and solid green lines respectively. Panel d) shows the thermodynamic trajectory, plotting order versus coarse-grained kinetic energy for the simulation evolution. the cut-off wavenumber $k_{c}$ is indicated with a dashed, grey line in panels b) and c), coinciding with spectral break-points, demonstrating the physical basis for the RFE operator (Eqs. 45).

### v4

@@ Abstract
To explain how dissipative active turbulence sustains inertial cascades, we construct a chiral model mapping overdamped agents to supersonic topological gas dynamics. Here, Mach cones function as acoustic horizons, shielding defect cores from the sound radiation of shallow water flows. We show that disorder activates a topological heat pump driving an inverse cascade toward a negative-temperature Onsager dipole, unless arrested into a vortex glass by insufficient dispersion. This mapping identifies a universality class unifying active swarms with the statistical mechanics of classical inviscid fluids.
@@ I Introduction
Active turbulence, the chaotic spatiotemporal dynamics observed in bacterial swarms [1, 2], cytoskeletal extracts [3], and synthetic colloidal flocks [4, 5], presents a challenge to statistical physics. Unlike classical inertial turbulence, which is driven by high Reynolds numbers ($\mathcal{R}e$), active turbulence arises in the overdamped limit ($\mathcal{R}e\approx 0$), driven by the continuous injection of energy at the microscale [6, 7].
Despite the viscous dominance in such systems, observations frequently reveal large-scale coherent structures reminiscent of inertial fluid flows [8]. This has ignited a debate regarding the universality class of active matter [3]. Current theories, largely built on the hydrodynamics of active nematics [9], predict steep energy spectra (typically $E(k)\sim k^{-4}$ or $k^{-3}$) dominated by the proliferation of half-integer defects [10, 11]. While such models successfully describe transient pattern formation, they generally fail to predict the robust inverse energy cascades ($k^{-5/3}$) observed in chiral active fluids [12] and 2D quantum turbulence [13].
Though active matter systems frequently and demonstrably exhibit viscous, dissipative turbulence [14, 15, 16], the observational discrepancy implies a missing link in the theoretical framework. While odd viscosity enables unidirectional flow [17, 18] and recent models suggest swarming can be entropy-driven [19], the thermodynamic end-state of such systems remains largely unexplored. Specifically, does the breaking of detailed balance [20] by intrinsic rotation sustain a chaotic steady state, or can it drive the system toward a specific macroscopic attractor?
In this letter, we propose that the answer to the conundrum lies in treating such systems as topological gases. By simulating chiral active matter as an ensemble of $N$ locally phase-coupled oscillators experiencing a Kuramoto-like phase-interaction [21], we demonstrate that active turbulence can exhibit a duality; a microscopic active bath dominated by enstrophy injection can simultaneously be described as a macroscopic renormalised fluid governed by the conservative advection of topological defects. By filtering out the microscopic singularities, we reveal that the simulated system undergoes a “hidden” inverse energy cascade, indistinguishable from Euler (frictionless, inviscid) turbulence. This cascade functions akin to a thermodynamic pump, driving the system toward an Onsager dipole, a negative-temperature condensate of vorticity [22], thereby unifying the phenomenology of biological swarms with the rigorous statistical mechanics of point vortices.
@@ II Methodology
The Model
We define the system as a chiral active fluid composed of $N=50,000$ particles, or discrete agents, in a bounded periodic domain $\mathcal{D}\in\mathbb{R}^{2}$. The state of the $i^{\text{th}}$ agent is defined by its position $\mathbf{r}_{i}(t)$ and its internal phase $\phi_{i}(t)\in[0,2\pi)$. The agents do not interact via pairwise collisions (as in granular matter) but through a local mean field. Following Ref. [21], we define the complex order parameter field $Z(\mathbf{r},t)$ as the convolution of the microscopic agent distribution with a finite-range interaction kernel $G$:
@@
where $G(|\mathbf{r}-\mathbf{r}^{\prime}|)$ is the Green’s function of the interaction (modeled as a Gaussian with kernel size $\sigma$), $R(\mathbf{r},t)$ is the local coherence (or order parameter, where $R\approx 1$ implies high local synchronisation and $R\approx 0$ implies a phase singularity, or defect core). $\Psi(\mathbf{r},t)$ is the local mean phase. The convolution kernel $G$ is interpreted in the context of chemical gradients or electric fields, which govern the dense active matter interactions (e.g., bacterial swarms). By modifying $G$, we can potentially simulate anisotropic media (like magnetized plasmas) without changing the agent logic.
@@
$\dot{\phi}_{i}=\underbrace{\omega_{i}}_{\text{Driver}}+\underbrace{a_{0}R(\mathbf{r}_{i})\sin(\Psi(\mathbf{r}_{i})-\phi_{i})}_{\text{Synchronisation force}}+\underbrace{\eta_{i}(t)}_{\text{Noise}},$ (2)
where $\omega_{i}$ are natural frequencies drawn from a power-law distribution $P(\omega)\sim\omega^{-n}$, $a_{0}$ is the coupling Strength, $\eta_{i}$ is a Gaussian white noise term (amplitude 0.15), and $n=1.5$ is a positive constant of order unity to facilitate the non-linear transfer of energy (see Refs. [21, 25]). $R_{i}$, the local order parameter, ensures that synchronisation is triggered by initially random local alignments in phase. $\omega_{i}$ (and the stochastic term $\eta_{i}$) prevent a “ferromagnetic collapse”, in which the system simply synchronises globally ($\phi_{i}=\text{const}$) causing a uniform ballistic drift. By forcing the particles to oscillate with an inherent frequency (or attempt to do so), the $\omega_{i}$ terms act as frustration, constantly injecting vorticity at the micro-scale, which, as we shall demonstrate, feeds an inverse energy cascade.
@@
where $v_{0}$ is the active swim speed and $\mathbf{r}_{i}=(x_{i},y_{i})$. The particles are therefore without conventional inertia: $\sum\mathbf{F}\neq m\mathbf{a}$; the system is fully overdamped, in a low Reynolds number limit [26]. The velocity field naturally forms vortices, introducing singularities, and so to define a valid continuum fluid, we introduce the renormalised fluid element (RFE), which we motivate theoretically in the Supplementary Materials.
At this point, We identify the characteristic scales in out model; the kernel size $\sigma$ and the active swim speed $v_{0}$, which yield the interaction time $\tau\equiv\sigma/v_{0}=3\text{ grid units}/0.5\text{ grid units/s}=6$ s, yielding the dimensionless units $\tilde{x}=x/\sigma$, $\tilde{t}=t/\tau$, and $\tilde{\omega}=\omega(\sigma/v_{0})$ (we subsequently drop the tilde). The system size becomes $\Lambda=2L/\sigma\approx 13$. A natural definition of the RFE is then afforded by $\tau$; we define the RFE as the center of mass of a particle’s trajectory during $\tau$, effectively course-graining the singularities:
@@
$r_{i}^{\text{RFE}}=\text{angle}\left(\frac{1}{\tau}\sum_{t^{\prime}=t-\tau}^{t}\hat{r}_{i}(t^{\prime})\right)\cdot\frac{\Lambda}{\pi},$ (4)
with $\hat{r}_{i}(t)$ being,
$\hat{r}_{i}(t)=\exp\left(i\cdot\frac{r_{i}(t)}{\Lambda}\cdot\pi\right),$ (5)
@@
being the complex phasor locations mapped to a unit circle (necessitated by the periodic boundary conditions). The transformation Eqs. (4, 5) constitute the regularization of the topological defects in $\phi_{i}$, which become physical flow vortices in $\mathbf{r}_{i}$.
@@
The correlation wavenumber of the local mean field is $k_{c}=2\pi/\sigma$, or simply $2\pi$ with normalized units, establishing a natural scale separation that marks the transition between the microscopic active bath ($k>k_{c}$) and the macroscopic hydrodynamic limit ($k<k_{c}$), with $k$ being the normalized wavenumber. The RFE operator effectively acts as a low-pass filter with cut-off $k_{c}$.
Figure 1: The emergence of the Onsager dipole in simulated chiral active matter, shown in seven temporal snapshots of the simulation space. The top row shows the RFE particle point-clouds $r_{i}^{\text{RFE}}$, the middle row shows topological charge $m$, and the bottom row shows inertial mass $R^{2}$, using the colorscales indicated.
Theoretical foundation
To bridge the discrete dynamics and the continuum fluid, we present, in the Supplementary Materials, a rigorous derivation proving that the dynamics produced by the model equations (Eqs. 15) are isomorphic to shallow water hydrodynamics with topography. A brief summary follows.
By decomposing the particle velocity into a mean flow with fluctuations, we find that the macroscopic velocity $\mathbf{u}$ is governed by the mean phase gradient $\nabla\Psi$ scaled by the local order parameter $R$: $\mathbf{u}\approx R\nabla\Psi$. While we derive a set of constraining equations that are isomorphic to shallow water hydrodynamics, there are striking differences: the effective inertial mass in our model scales with synchronization:
$\lambda=R^{2}.$ (6)
At the same time, the effective pressure scales with disorder:
$P_{\text{eff}}=\frac{v_{0}^{2}}{2}(1-R^{4}).$ (7)
Crucially, a stiff active pressure enforces a sound speed proportional to order ($c_{s}\propto R$). Since the inertial velocity also scales with $R$, this locks the flow into an inherently supersonic state (Mach number >1>1). Consequently, defect cores ($R\rightarrow 0$) are isolated behind Mach cones, creating sonic horizons that prevent radiative decay. While fast compressible shocks disperse inertial energy locally, the time-averaged background flow remains incompressible. This effectively traps energy within the topological degrees of freedom, meaning the macroscopic dynamics are governed strictly by the conservation of topological charge.
@@
As the mean field $\Psi(r)$ is the argument of a single-valued complex field (Eq. 1), its circulation along any closed loop $\partial\Sigma$ enclosing a defect core is necessarily quantized [27, 28], mirroring the winding number stability in discrete oscillator ensembles [29, 21]:
@@
$\oint_{\partial\Sigma}\nabla\Psi\cdot d\mathbf{l}=2\pi m.\quad(m\in\mathbb{Z})$ (8)
Thus, the interaction energy of the ensemble is given by the Kirchhoff-Onsager Hamiltonian [30, 31]:
@@
This Hamiltonian confirms that while individual agents are overdamped, the renormalised fluid behaves as an inviscid point-vortex gas of topological defects on long timescales.
The dynamic Onsager attractor
In a bounded domain $\mathcal{D}$, the phase space volume $\Omega(E)$ eventually decreases as energy increases, implying a negative temperature regime ($T<0$) where entropy favors the merging of like-signed vortices [22]. However, in our active chiral matter model, the dipole is continuously eroded by supersonic shocks which thermalise inertial energy back into the active bath. Consequently, the Onsager dipole emerges as a dynamic attractor, where an active bath constantly rebuilds topological order using newly thermalised energy, sustaining the system in the negative-temperature Onsager regime. Crucially, the mechanism parallels observed inverse-energy cascades [1], motility-induced phase separation [32], and entropy-driven ordering [19] in active matter studies.
Spectral scaling
@@
Figure 2: Energy characteristics of the simulation run in Figure 1. Panel a) shows the defect (phase singularity) count, with a $t^{-0.75}$ powerlaw scaling [38] shown with a dashed, green line. Panel b) shows the spectral flux (Eq. 10) for the RFE (blue, left $y$-axis) and raw (red, right $y$-axis) fields, taking the median of the entire simulation. Panel c) shows the energy spectrum $E(k)$ for the RFE (blue) and raw (red) fields; a spectral scaling of $k^{-5/3}$ and $k^{-8/3}\approx k^{-2.67}$ is indicated with dotted and solid green lines respectively, and $k_{c}$ is indicated with a dashed, grey line, as in Panel b). Panel d) shows the thermodynamic trajectory, plotting order versus coarse-grained kinetic energy for the simulation evolution.
@@
which sums contributions of all discrete wave vectors $\mathbf{q}=(q_{x},q_{y})$ that fall within the annular shell where the magnitude $|\mathbf{q}|\approx k$, to the real part of the Fourier transform of the advective transport term, where $\hat{\mathbf{v}}_{\mathbf{q}}^{*}$ is the complex conjugate of the velocity field in Fourier space at wave vector $\mathbf{q}$. A negative plateau in $\Pi(k)$ will conclusively demonstrate an inverse cascade.